On January 1, 2018, Diana Corporation (a lessor) signed to lease a machinery for eight-year. The lease was noncancellable. The terms of the lease called for Diane to make annual payments of $220,000 starting January 1, 2019 for eight years with the title reverting to the Diana at the end of the lease term. The contract stated that Diane guaranteed the residual value of the machine is $10,000 upon end of lease term. The machinery originally costed Diana $750,000, has an estimated useful life of 15 years and no salvage value. Both Diana and Diane use the straight-line method of depreciation for all of its fixed assets. Diana used an effective interest rate of 8% on this lease, and the rate is known by Diane.

(PV ordinary annuity (8%, t=8) = 5.74664; FV ordinary annuity (8%, t=8) = 10.63663; PV annuity due (8%, t=8) = 6.2064, PV single sum (8%, t=8) = 0.5403).

On January 1, 2018, Diana (lessor) recorded lease receivable of _______ and during 2018, Diana (lessor) recorded depreciation expense of ________ regarding this asset.
Answer #1

Lease receivable = (annual payments*present value of annuity due of n=8 i = 8%)+ (residual value* present value of n= 8, i = 8%) (220000*6.2064)+(10000*0.5403) = $1370811

Depreciation expense = cost /estimated life = 750000/15 = $50000